Allowance Method for Accounting for Bad Debts

At the beginning of 2017, EZ Tech Company's Accounts Receivable balance was $147,000, and the balance in Allowance for Doubtful Accounts was $2,500. EZ Tech's sales in 2017 were $1,100,000, 70% of which were on credit. Collections on account during the year were $710,000. The company wrote off $4,000 of uncollectible accounts during the year.

Required:

1. Identify and analyze the sales during 2017.

Activity
Accounts
Statement(s)

How does this entry affect the accounting equation?
If a financial statement item is not affected, select "No Entry" and leave the amount box blank. If the effect on a financial statement item is negative, i.e, a decrease, be sure to enter the answer with a minus sign.

Balance Sheet Income Statement
Stockholders' Net
Assets = Liabilities + Equity Revenues Expenses = Income

Identify and analyze the transactions related to the collections of cash during 2017.

Activity
Accounts
Statement(s)

How does this entry affect the accounting equation?
If a financial statement item is not affected, select "No Entry" and leave the amount box blank. If the effect on a financial statement item is negative, i.e, a decrease, be sure to enter the answer with a minus sign.

Balance Sheet Income Statement
Stockholders' Net
Assets = Liabilities + Equity Revenues Expenses = Income

Identify and analyze the transactions related to the write-offs of accounts receivable during 2017.

Activity
Accounts
Statement(s)

How does this entry affect the accounting equation?
If a financial statement item is not affected, select "No Entry" and leave the amount box blank. If the effect on a financial statement item is negative, i.e, a decrease, be sure to enter the answer with a minus sign. Remember: if a contra account is increased, it will have the effect of decreasing the corresponding financial statement item. If a contra account is decreased, it will have the effect of increasing the corresponding financial statement item.

Balance Sheet Income Statement
Stockholders' Net
Assets = Liabilities + Equity Revenues Expenses = Income

2. Identify and analyze the adjustments to recognize bad debts assuming that (a) bad debts expense is 2% of credit sales and (b) amounts expected to be uncollectible are 7% of the year-end accounts receivable.

a. Identify and analyze the adjustments to recognize bad debts assuming that bad debts expense is 2% of credit sales.

Activity
Accounts
Statement(s)

How does this entry affect the accounting equation?
If a financial statement item is not affected, select "No Entry" and leave the amount box blank. If the effect on a financial statement item is negative, i.e, a decrease, be sure to enter the answer with a minus sign. Remember: if a contra account is increased, it will have the effect of decreasing the corresponding financial statement item. If a contra account is decreased, it will have the effect of increasing the corresponding financial statement item.

Balance Sheet Income Statement
Stockholders' Net
Assets = Liabilities + Equity Revenues Expenses = Income

b. Identify and analyze the adjustments to recognize bad debts assuming that amounts expected to be uncollectible are 7% of the year-end accounts receivable.

Activity
Accounts
Statement(s)

How does this entry affect the accounting equation?
If a financial statement item is not affected, select "No Entry" and leave the amount box blank. If the effect on a financial statement item is negative, i.e, a decrease, be sure to enter the answer with a minus sign. Remember: if a contra account is increased, it will have the effect of decreasing the corresponding financial statement item. If a contra account is decreased, it will have the effect of increasing the corresponding financial statement item.

Balance Sheet Income Statement
Stockholders' Net
Assets = Liabilities + Equity Revenues Expenses = Income

3. What is the net realizable value of accounts receivable on December 31, 2017, under each assumption in part (2)?

Using the percentage of sales approach, the net realizable value of the receivables is?
$

Using the percentage of year-end receivables approach, the net realizable value of the receivables is?
$

4. The recognition of bad debts expense_________ the net realizable value by the amount recorded in bad debts expense and the_________.The write-off of accounts__________ the net realizable value.
Answer #1
1 Sales:
Activity Operating
Accounts Cash increase,Accounts receivable increase,Sales increases
Statement(s) Income statement,balance sheet and cash flow statement
Balance sheet Income statement
Assets = Liabilities + Stockholder's Equity Revenues - Expenses = Net income
Cash (1100000*30%) 330000 No entry 0 1100000 Sales 1100000 No entry 0 1100000
Accounts receivable
(1100000*70%) 770000
Collections of cash during 2017:
Activity Operating
Accounts Cash increase,Accounts receivable decrease
Statement(s) Balance sheet only
Balance sheet Income statement
Assets = Liabilities + Stockholder's Equity Revenues - Expenses = Net income
Cash 710000 No entry 0 0 No entry 0 No entry 0 0
Accounts receivable -710000
Write off of uncollectible accounts:
Activity NA
Accounts Allowance for doubtful accounts decrease,Accounts receivable decrease
Statement(s) Balance sheet only
Balance sheet Income statement
Assets = Liabilities + Stockholder's Equity Revenues - Expenses = Net income
Allowance for doubtful accounts 4000 No entry 0 0 No entry 0 No entry 0 0
Accounts receivable -4000
2
(a) Bad debt expense=Credit sales*2%=770000*2%=$ 15400
Activity NA
Accounts Bad debt expense increase,Allowance for doubtful accounts increase
Statement(s) Balance sheet only
Balance sheet Income statement
Assets = Liabilities + Stockholder's Equity Revenues - Expenses = Net income
Allowance for doubtful accounts -15400 No entry 0 -15400 No entry 0 Bad debt expense 15400 -15400
(b) Desired balance in allowance for doubtful accounts=Year-end accounts receivable*7%
Year-end accounts receivable:
$
Beginning balance 147000
Add: Credit sales 770000
917000
Less:Collections 710000
Ending balance 207000
Desired balance in allowance for doubtful accounts=207000*7%=$ 14490
Current balance in allowance for doubtful accounts=2500-4000=-1500=$ 1500 debit
Bad debt expense to be recognized=14490+1500=$ 15990
Activity NA
Accounts Bad debt expense increase,Allowance for doubtful accounts increase
Statement(s) Balance sheet only
Balance sheet Income statement
Assets = Liabilities + Stockholder's Equity Revenues - Expenses = Net income
Allowance for doubtful accounts -15990 No entry 0 -15990 No entry 0 Bad debt expense 15990 -15990

3 Using the percentage of sales approach, the net realizable value of the receivables is: Net realizable value=Year end accou